OpenStudy (anonymous):

sin = 9/16 find the missing value

OpenStudy (mathmale):

Kito, you haven't included a variable in your equation. Could you do that, please? Please re-read the original question and make certain you copy it down in is entirety.

OpenStudy (anonymous):

what the question says is find the missing value sin=9/16 I got the answer 34.23 but there was no picture or anything. my problem is I don't know the steps to get the answer it also said round to the nearest hundredth

OpenStudy (mathmale):

Kito: What I was driving at is that each trig function must have an argument (input variable). sin cannot stand alone. What you meant was sin x = 9/16. Here you are told what the value of the sine of the angle x is (9/16) and are asked to find the value of the angle, x. Have you seen and/or heard of the "inverse sine" function before?

OpenStudy (anonymous):

no I just started this chapter and there was 20 questions and only got 5 right also my teacher told me to down load the trig calculator and it was no help

OpenStudy (anonymous):

just looked up inverse sine function ??????????????? do you know of a web site that might show step by step for dummies:) because I am sooooo lost. thanks for your help!!!!!

OpenStudy (anonymous):

if sin(A) = (9/16) A = arcsin(9/16), if that is what you want to find

OpenStudy (anonymous):

You can draw a triangle too if you want to visualize it

OpenStudy (anonymous):

im not sure so I know the answer is 34.23 but what I don't know is how to go about getting that answer

OpenStudy (anonymous):

okay so I read that you take the opposite and the hypotenuse and divide them but I got something very different

OpenStudy (mathmale):

Here's the basic idea: IF sin x = y, then \[\sin^{-1} y=x\]

OpenStudy (anonymous):

*or mathmale can do it if that's what you are already doing

OpenStudy (anonymous):

hey im loving you both right now

OpenStudy (mathmale):

Example: sin (30 deg) = 1/2 and\[\sin^{-1} \frac{ 1 }{ 2 }=30~\deg~ or~ \pi/6~~ radians\]

OpenStudy (mathmale):

So, if you are told that sin x = 9/16, how would you obtain x? Hint: use the inverse function discussed immediately above. Larry: Care to set this up?

OpenStudy (anonymous):

sin 30 = 9/16????

OpenStudy (anonymous):

sin 30 DEGREES is (1/2)

OpenStudy (mathmale):

actually, it'd be sin x = 9/16. Larry will probably tell you next to take the inverse sine of both sides of this equation sin x = 9/16. Larry?

OpenStudy (mathmale):

Important introduction (or review) to inverse trig functions: \[\sin(\sin ^{-1}x)=x\]

OpenStudy (mathmale):

and\[\sin ^{-1}\sin x=x\]

OpenStudy (anonymous):

You have to be careful when putting down numbers. If the degrees isn't there, I was told years ago (although it might be different now), that it is implied to be in radians. We both already stated thing about acrsin (or opposite sin, not sin to the power of -1 )...and I was going to make a drawing but failed at it horribly, but yeah, look at what mathmale showed

OpenStudy (mathmale):

We say, "inverse functions undo each other." Given x, if you find sin x, and then want your x back, use the inverse sine function to get it back.

OpenStudy (anonymous):

Just how x+1-1 = x

OpenStudy (anonymous):

*well, kind of

OpenStudy (anonymous):

thank you both so very much u took the time to explain it helps a ton.

OpenStudy (mathmale):

So, in summary, if sin x = 9/16, x = arcsin 9/16 = approx. 34.23 degrees, using a TI-83 Plus calculator.